OpenStudy (anonymous):

How do I find the domain and range of these functions? f(x)= sinx - 1 g(x)= cosx - 1

OpenStudy (anonymous):

u knw the domain and range of their parent functions : sinx, cosx ?

OpenStudy (anonymous):

-1 <= sinx <= 1 and -1 <= cosx <= 1?

OpenStudy (anonymous):

yes, those are the ranges

OpenStudy (anonymous):

range of sinx :- -1 <= sinx <= 1 subtract 1 through out -1-1 <= sinx-1 <= 1-1 -2 <= sinx-1 <= 0 so, range of sinx-1 is : [-2, 0]

OpenStudy (anonymous):

why did you substract one?

OpenStudy (anonymous):

cuz we're after finding range of sinx-1

OpenStudy (anonymous):

and we knw the range of sinx

OpenStudy (anonymous):

so subtracting 1 gives u range of sinx-1

OpenStudy (anonymous):

What abotu cosx?

OpenStudy (anonymous):

do it same way

OpenStudy (anonymous):

But on my book it says that the range of sinx-1 is : [-2, 0] and that the range of cosx-1 is R... Why is that?

OpenStudy (anonymous):

range of sinx-1 and cosx-1 is : [-2, 0] domain of sinx-1 and cosx-1 is : R

OpenStudy (anonymous):

Ok not I get it

OpenStudy (anonymous):

Can I ask you one more thing?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Knowing the domain and the range of these functions, how to do I find a function composition that is f(g(x))?

OpenStudy (anonymous):

easy, f( g(x) ) domain = range of g(x)

OpenStudy (anonymous):

so, domain of f(g(x)) = [-2, 0]

OpenStudy (anonymous):

and the function? how can I find it?

OpenStudy (anonymous):

oh you want to find function ?

OpenStudy (anonymous):

yea, I'm sorry if my english is not good

OpenStudy (anonymous):

f(x)= sinx - 1 g(x)= cosx - 1 f(g(x)) = f(cosx-1) = sin(cosx-1) - 1

OpenStudy (anonymous):

thats teh function

OpenStudy (anonymous):

Ok thank you!

OpenStudy (anonymous):

u wlc (:

OpenStudy (anonymous):

HEY dear that is easy you have to know that cos and sin can not exceed -1 and 1 so check between this range for more write for me on www.facebook.com/beastargroup